Hello, I would like help on the steps to put the two points ( 4,1) and ( -1,7) in ax + by = c form. ( standard form)
Please and thank you.

Answer by Alan3354(69443)

Here's an example that's similar.
================================
I need find an equation of the line in standard form passing through two lines.
(5/4,3/4) and (2/3,1/3)
----------
Find the slope, m
m = diffy/diffx = (5/12)/(3/12)
m = 5/3
------
Use y = mx + b and either point to find b, the y-intercept
3/4 = (5/3)*(5/4) + b
9/12 = 15/12 + b
b = -6/12
b = -1/2
-----------
--> y = (5/3)x - 1/2
or
10x - 6y = 3